# Prediction and identification of point defect fingerprints in X-ray photoelectron spectra of TiNx with 1.18 1.37

### v2

@@ 3.5 Ti vacancy quantification
Based on DFT calculations, the observed changes in the N 1s spectra with increasing off-stoichiometry can be explained by Ti vacancies, N tetrahedral interstitials or Ti interstitials. The formation of Ti interstitials can be ruled out, since they are predicted to exhibit a distinct fingerprints also in the Ti 2p spectra, with a well separated peak at lower binding energies, shifted by 1.5–2.3 eV with respect to the main TiN peak, and no such peak is observed in the here studied films as shown in Figure 3b). This is expected as the Ti interstitials have a high formation energies and are furthermore not favored by the N-rich film stoichiometry. On the contrary, the presence of both Ti vacancies and N tetrahedral interstitials is consistent with the film composition measured by the ion beam analysis (Table 1) and XRD (Figure 1) results. Only the tetrahedral N interstitials are predicted to be detectable in the XPS spectra. The split N interstitials are more stable than the tetrahedral ones [55], however, as HPPMS deposition is a highly energetic process, creating N tetrahedral interstitials cannot be ruled out. Because the N tetrahedral interstitials have a fingerprint that is indistinguishable from the fingerprint of the Ti vacancies, while the split N interstitials do not have a signal overlapping with the signal of Ti vacancies, it is important to estimate the respective N interstitial populations.
@@
where is energy, is peak position, is peak broadening parameter (its value is close to full width at half the maximum intensity), and stands for fraction of the Lorentzian part in % of a peak were employed. The GL(70) lineshape was used for the main TiN-like (0V) peak, while GL(30) line shapes were utilized for the 1V and 2V peaks and GL(0) for the satellite peak. Peak intensities, positions, and broadening were free-fitting parameters, however, only a single shared broadening parameter was fitted for the 1V and 2V peaks to reduce fit ambiguity. A Tougaard-like background shape was utilized during the fitting. The fits of the N 1s regions of all samples are shown in Figure 5, while the fit parameters are summarized in Table 3.5. Using this four-peak model, it was possible to achieve near-perfect fits for all the studied samples. Additionally, the fit parameters are very stable between the fits and the fitted peak positions are in good agreement with the predicted shifts. Specifically, the positions of the three main peaks are in the range of 397.12–397.19 eV, 396.49–396.52 eV, and 396.91–396.07 eV for the 0V, 1V, and 2V components, respectively. This is in excellent agreement with the here calculated value of eV for the shift between 0V and 1V peak as well as the predicted and experimentally observed shifts between corresponding components in TiON [29].
The fitting was followed by a simplified version of the analysis of the relative peak ratios from [29], enabled by the fact that there is no oxidation of the TiN layer due to the Al capping, and, therefore, no need for corrections compensating the TiN oxidation. Vacancy concentration can be directly quantified from the three N 1s main components (0V, 1V, 2V) ratios. Assuming a cubic TiN monocrystal, i.e., neglecting the grain boundaries, with random vacancy distribution, the probability of finding an N atom with Ti first neighbors, or equivalently with neighboring Ti vacancies, is [29]
@@
During the vacancy quantification, we search for a which minimizes the squared differences between the experimental (fitted) components and the theoretical ones, where both the experimental and theoretical functions were previously renormalized to sum to 1. Only the first three components are considered, i.e., corresponding to the N atoms with 0, 1, and 2 Ti vacancies in the first coordination shell. The normalization is needed for the experimental components, as opposed to using directly the fractions from Table 3.5, since the fit also includes the satellite peak. For the theoretical probabilities, the normalization is equivalent to neglecting the components corresponding to the N atoms with more than two vacancies in the first coordination shell. There is no experimental evidence for a corresponding signal in our data and the probability of finding such configuration at small is very low; therefore, this is a reasonable simplification. This approach is consistent with that employed in Ref. [29], but it is significantly more robust due to the capping layer, eliminating the need for additional parameters to account for oxidation, which in turn reduces uncertainty.
@@
The theoretical are shown in Figure 6 together with the fitted ratios for the four here studied samples. The axis positions of the experimental points correspond to the fitted using the aforementioned procedure and is thus the determined Ti vacancy concentration on the metal sublattice.
@@ 4 Conclusions
Based on ab initio predictions, it is shown that several point defects in TiN, specifically Ti vacancies, N tetrahedral interstitials and Ti interstitials, induce N 1s and/or Ti 2p core electron binding energy shifts in the first N and/or Ti neighbors, with a magnitude large enough to be detectable with standard laboratory X-ray photoelectron spectroscopy. The here presented calculations predict that significant N 1s BE shifts are associated with N tetrahedral interstitials and Ti vacancies, with N 1s shifts of -0.54 eV for first N neighbors of the Ti vacancy and N 1s shifts of -0.54 eV for first N neighbors of the tetrahedral N interstitial. However, while is is not possible to distinguish the Ti vacancies from N tetrahedral interstitials by the evaluation of N 1s binding energy shifts, we have also shown that the N tetrahedral interstitial is not stable at the room temperature. It quickly transform to the lower energy split interstitials and the population of tetrahedral N interstitials in TiN is negligible as a result. Therefore, we have shown that the N 1s signal at lower binding energies can be unambiguously attributed to Ti vacancies, Ti vacancies can be unambiguously detected in the XPS measurements of Al-capped TiN and the recently published XPS Ti vacancy quantification method for Ti1-δOxN1-x thin films can be adapted also for TiN films. This was successfully experimentally verified by XPS measurements of Al-capped TiNx thin films with 1.18(5) 1.37(5), where the Ti vacancy concentration on the metal sublattice ranged from % to %, respectively. The largest predicted Ti 2p BE shifts are associated with Ti interstitials, however no evidence for the formation of Ti interstitials was detected in the as deposited thin films. Hence, N rich TiN accommodates off-stoichiometry by both, formation of Ti vacancies as well as N split interstitials. Our results highlight that the ab initio-guided XPS-based point defect characterization method is suitable for the detection and quantification of several point defects in Al-capped TiNx.
@@ Acknowledgments
This research was funded by German Research Foundation (DFG, SFB-TR 87/3) ”Pulsed high power plasmas for the synthesis of nanostructured functional layer”. This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254) and project LM2023039. The authors also gratefully acknowledge the computing time granted by the JARA Vergabegremium and provided on the JARA Partition part of the supercomputer CLAIX at RWTH Aachen University (project JARA0151) and computational resources provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic. The NEB calculations have been achieved using the Vienna Scientific Cluster (VSC). Accelerator operation at Uppsala University has been supported by the Swedish research council VR-RFI (#2019_00191).

### v3

@@ 3.5 Ti vacancy quantification
Based on DFT calculations, the observed changes in the N 1s spectra with increasing off-stoichiometry can be explained by Ti vacancies, N tetrahedral interstitials or Ti interstitials. The formation of Ti interstitials can be ruled out, as the Ti interstitials have a high formation energies and are furthermore not favored by the N-rich film stoichiometry. On the contrary, the presence of both Ti vacancies and N tetrahedral interstitials is consistent with the film composition measured by the ion beam analysis (Table 1) and XRD (Figure 1) results. Only the tetrahedral N interstitials are predicted to be detectable in the XPS spectra. The split N interstitials are more stable than the tetrahedral ones [56], however, as HPPMS deposition is a highly energetic process, creating N tetrahedral interstitials cannot be ruled out. Because the N tetrahedral interstitials have a fingerprint that is indistinguishable from the fingerprint of the Ti vacancies, while the split N interstitials do not have a signal overlapping with the signal of Ti vacancies, it is important to estimate the respective N interstitial populations.
@@
where is energy, is peak position, is peak broadening parameter (its value is close to full width at half the maximum intensity), and stands for fraction of the Lorentzian part in % of a peak were employed. The GL(70) lineshape was used for the main TiN-like (0V) peak, while GL(30) line shapes were utilized for the 1V and 2V peaks and GL(0) for the satellite peak. Peak intensities, positions, and broadening were free-fitting parameters, however, only a single shared broadening parameter was fitted for the 1V and 2V peaks to reduce fit ambiguity. A Tougaard-like background shape was utilized during the fitting. The fits of the N 1s regions of all samples are shown in Figure 5, while the fit parameters are summarized in Table 3. Using this four-peak model, it was possible to achieve near-perfect fits for all the studied samples. Additionally, the fit parameters are very stable between the fits and the fitted peak positions are in good agreement with the predicted shifts. Specifically, the positions of the three main peaks are in the range of 397.12–397.19 eV, 396.49–396.52 eV, and 395.91–396.07 eV for the 0V, 1V, and 2V components, respectively. The 1V and 2V components exhibit negative shifts of 0.62–0.67 eV and 1.12–1.21 eV, respectively. These ranges reflect the variability among the four fits of different samples, relative to 0V (TiN) component. This is in excellent agreement with the here predicted value of eV for the shift between 0V and 1V peak as well as the predicted and experimentally observed shifts between corresponding components in TiON [29].
The fitting was followed by a simplified version of the analysis of the relative peak ratios from [29], enabled by the fact that there is no oxidation of the TiN layer due to the Al capping, and, therefore, no need for corrections compensating the TiN oxidation. The simplified analysis comes at the expense of more complex deposition procedure and can limit the applicability of this method in some cases, e.g., if the Al deposition step can not be easily added to the existing deposition workflow. Vacancy concentration can be directly quantified from the three N 1s main components (0V, 1V, 2V) ratios. Assuming a cubic TiN monocrystal, i.e., neglecting the grain boundaries, with random vacancy distribution, the probability of finding an N atom with Ti first neighbors, or equivalently with neighboring Ti vacancies, is [29]
@@
During the vacancy quantification, we search for a which minimizes the squared differences between the experimental (fitted) components and the theoretical ones, where both the experimental and theoretical functions were previously renormalized to sum to 1. Only the first three components are considered, i.e., corresponding to the N atoms with 0, 1, and 2 Ti vacancies in the first coordination shell. The normalization is needed for the experimental components, as opposed to using directly the fractions from Table 3, since the fit also includes the satellite peak. For the theoretical probabilities, the normalization is equivalent to neglecting the components corresponding to the N atoms with more than two vacancies in the first coordination shell. There is no experimental evidence for a corresponding signal in our data and the probability of finding such configuration at small is very low; therefore, this is a reasonable simplification. This approach is consistent with that employed in Ref. [29], but it is significantly more robust due to the capping layer, eliminating the need for additional parameters to account for oxidation, which in turn reduces uncertainty.
@@
The theoretical are shown in Figure 6 together with the fitted ratios for the four here studied samples. The axis positions of the experimental points correspond to the fitted using the aforementioned procedure and is thus the determined Ti vacancy concentration on the metal sublattice. The determined vacancy concentrations are , , , and percent of Ti vacancies on the metal sublattice for the TiN1.18, TiN1.23, TiN1.35 and TiN1.37, respectively. The vacancy concentrations are lower than would be expected based on the ERDA/RBS film stoichiometry under the assumption of Ti vacancies only, and the most likely explanation for this discrepancy is presence of N interstitials in the films. The insensitivity of the XPS-based point defect detection method to the presence of N interstitial therefore can be considered as a strength of the method when determining the Ti vacancies, as opposed to possible estimates from the stoichiometry or a stress-free lattice parameter, since the stoichiometry or lattice parameter measurements are always influenced by the sum of effects from both N interstitials and Ti vacancies with no option to distinguish the individual contributions. It most be noted that in some cases (most notably with TiN1.37), that the experimentally determined 0V, 1V and 2V component ratios are not a perfect fit for the theoretically-modeled component ratios. Specifically, the experimental 2V component ratio would better correspond to higher value, while the experimental 0V and 1V ratios would suggest a lower value. As noted before, our model in Eq. (3) assumes a random vacancy distribution. If the vacancy distribution is not completely random, some disagreement between the theoretical and experimental component ratios would be expected. It would be therefore theoretically possible to detect not only the vacancy concentration but also provide some information about their distribution, however, such a theoretical model is beyond the scope of this work.
@@ 4 Conclusions
Based on ab initio predictions, it is shown that several point defects in TiN, specifically Ti vacancies and N tetrahedral interstitials induce N 1s core electron binding energy shifts in the first N neighbors, with a magnitude large enough to be detectable with standard laboratory X-ray photoelectron spectroscopy. The here presented calculations predict that significant N 1s BE shifts are associated with N tetrahedral interstitials and Ti vacancies, with N 1s shifts of -0.54 eV for first N neighbors of the Ti vacancy and N 1s shifts of -0.54 eV for first N neighbors of the tetrahedral N interstitial. However, while it is not possible to distinguish the Ti vacancies from N tetrahedral interstitials by the evaluation of N 1s binding energy shifts, we have also shown that the N tetrahedral interstitial is not stable at the room temperature. It quickly transform to the lower energy split interstitials and the population of tetrahedral N interstitials in TiN is negligible as a result. Therefore, we have shown that the N 1s signal at lower binding energies can be unambiguously attributed to Ti vacancies, Ti vacancies can be unambiguously detected in the XPS measurements of Al-capped TiN and the recently published XPS Ti vacancy quantification method for Ti1-δOxN1-x thin films can be adapted also for TiN films. By showing that the N 1s BE shift in the vicinity of a metal vacancy is independent of the presence of O in the film, we also demonstrate the potential for application to other NaCl-structured nitrides. The proposed method was successfully experimentally verified by XPS measurements of Al-capped TiNx thin films with 1.18(5) 1.37(5), where the Ti vacancy concentration on the metal sublattice ranged from % to %, respectively. Importantly, our application of the Al capping simplifies significantly the previously proposed vacancy quantification methodology, leads to more robust results and we have also presented a rigorous capping quality validation method based on the angle-resolved XPS. Hence, N rich TiN accommodates off-stoichiometry by both, formation of Ti vacancies as well as N split interstitials. Our results highlight that the ab initio-guided XPS-based point defect characterization method is suitable for the detection and quantification of several point defects in Al-capped TiNx.
@@ Acknowledgments
This research was funded by German Research Foundation (DFG, SFB-TR 87/3) ”Pulsed high power plasmas for the synthesis of nanostructured functional layer”. This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the project LM2023039. The authors also gratefully acknowledge the computing time granted by the JARA Vergabegremium and provided on the JARA Partition part of the supercomputer CLAIX at RWTH Aachen University (project JARA0151) and computational resources provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic. The NEB calculations have been achieved using the Vienna Scientific Cluster (VSC). Accelerator operation at Uppsala University has been supported by the Swedish research council VR-RFI (#2019_00191).
Data availability
Ab initio simulations are available under the CC BY 4.0 licence in the NOMAD archive [67]. Experimental data will be made available on request.